OpenStudy (anonymous):

Simplify each expression: 1) -6(a+8) 2) 4(1+9x) 3) 6(-5n+7) 4) (9m+10)x2

hartnn (hartnn):

use distributive property a(b+c) =ab+ac so what will be -6(a+8) = ?

OpenStudy (anonymous):

i still dont get it.

hartnn (hartnn):

 what is -6*8 ?